OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is 5x - 3 and whose width is x + 7? 5x2 - 21 5x2 + 9x - 21 5x2 + 32x - 21 5x2 - 32x - 21

OpenStudy (anonymous):

Expand$$ (5x-3)(x+7) = 5x^2 + 32x - 21$$

OpenStudy (anonymous):

Area = length * width \[(5x-3)(x+7)\]FOIL it \[5x ^{2}+35x-3x-21\] \[5x ^{2}-32x-21\]

OpenStudy (anonymous):

no this is right (5x−3)(x+7)=5x2+32x−21

OpenStudy (anonymous):

@deezil_wood: Check for the sign's ;)

OpenStudy (anonymous):

oh yeah duh don't know where that -32 came from lol

OpenStudy (anonymous):

$$35x-3x = +32x \neq -32x $$

OpenStudy (anonymous):

it's happens, obvious typo :)

OpenStudy (anonymous):

Yes thx for the explanation :p

OpenStudy (anonymous):

you are welcome :P :D